# Optimal Attention Temperature Enhances In-Context Learning under Distribution Shift

### v1

@@ Appendix F Derivation of the pretraining for ICL by mimicking the Bayes-optimal estimator
Here, we derive the pretraining of the linearized attention model by mimicking the Bayes-optimal ridge estimator (9). Recall that the prediction of the linearized attention model is
@@
which makes the prediction of the linearized attention model approximately equal to that of the Bayes-optimal regression. Furthermore, we consider $\tau=1$ for the pretraining. Let’s first focus on $\hat{{\bm{w}}}_{Att}({\bm{C}}_{xx},{\bm{C}}_{xy},C_{yy};{\bm{M}},{\bm{V}})$ as follows
@@
as derived in Appendix A. Looking at equations (53) and (52) together, we can see that setting the parameters as follows would make $\hat{{\bm{w}}}_{Att}=\hat{{\bm{w}}}_{Bayes}$ hold
@@ Appendix G Characterization of generalization error for ICL under distribution shift
by the law of total expectation since ${\bm{w}}_{diff}$ is independent of ${\bm{x}}_{l}$. Note that when writing (65), we safely ignore terms with $(1/l)\bar{{\bm{x}}}_{l}\bar{{\bm{x}}}_{l}^{T}{\bm{v}}_{21}$ in (63) since they vanish as $l\to\infty$ by Assumptions 3.1-3.2 and 4.5. Letting ${\bm{A}}:={\bm{\mu}}_{x}{\bm{\mu}}_{x}^{T}+{\bm{\Sigma}}_{x}$, we write
@@
where we first use the independence of the random variables and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ to simplify the equation. Then, we apply Lemma G.1 and use the fact that $\mathbb{E}[\bar{{\bm{x}}}_{i}\bar{{\bm{x}}}_{i}^{T}]={\bm{\Sigma}}_{x}$ to get the penultimate line. Finally, we drop the vanishing terms and simplify the result using Assumptions 3.1-3.2 and 4.5 in order to reach the last line.
@@
where we again first use the independence of the random variables and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. Then, we apply basic algebraic manipulations. To reach the penultimate line, we utilize Lemma G.1 together with the fact that $\mathbb{E}[\bar{{\bm{x}}}_{i}\bar{{\bm{x}}}_{i}^{T}]={\bm{\Sigma}}_{x}$. Using Assumptions 3.1-3.2 and 4.5, we reach the last line.
@@ Appendix I Analysis of optimal temperature for ICL under distribution shift
$\tau_{optimal}$ $:=\operatorname*{arg\,min}_{\tau}\mathcal{G}(\tau;{\bm{V}},{\bm{M}}),$ (115)
@@
Since $a,b>0$, we reach $\mathcal{G}^{\prime\prime}\left(\frac{2a}{b}\right)=\frac{b^{4}}{8a^{3}}>0$, which means the function has a minimum at $\tau=\frac{2a}{b}$. Therefore, $\tau_{optimal}=\frac{2a}{b}$ is the solution minimizing the generalization error $\mathcal{G}(\tau)$. Writing $a,b$ back into the optimal solution, we get
$\tau_{optimal}=\frac{2\text{Tr}\left({\bm{A}}{\bm{M}}_{11}^{T}{\bm{F}}_{1}{\bm{M}}_{11}\right)}{\text{Tr}\left({\bm{A}}\left({\bm{F}}_{2}{\bm{M}}_{11}+{\bm{M}}_{11}^{T}{\bm{F}}_{2}^{T}\right)\right)},$ (122)
which concludes our derivation of the optimal temperature $\tau_{optimal}$.
@@
In this section, we extract a mathematical heuristic from the optimal temperature in Theorem 4.7 that can be applied to ICL settings beyond our existing setting involving linearized attention and regression tasks. Specifically, we consider Transformers employing standard softmax attention. Recall that the attention temperature scales the pre-softmax scores (i.e., $({\bm{K}}{\bm{Z}})^{\top}({\bm{Q}}{\bm{Z}})$ in (1)), thereby controlling the variance of the final scores. Since the optimal temperature depends on the distribution of these scores, it can be naturally characterized by the moments of that distribution. Our central intuition is that the optimal temperature identified in Theorem 4.7 relates directly to the first two moments of the pre-softmax scores. Although this optimal temperature was derived for linearized softmax attention, the insight remains relevant for softmax attention because the two mechanisms behave similarly in the regime considered (see Appendix D).
@@
$\tau_{\mathrm{optimal}}=\frac{v_{22}\operatorname{Tr}\!\left({\bm{\Sigma}}_{x}{\bm{M}}_{11}{\bm{\Sigma}}_{x}{\bm{M}}_{11}^{\top}{\bm{\Sigma}}_{x}\right)}{\frac{1}{2}\operatorname{Tr}\!\left({\bm{\Sigma}}_{x}\left({\bm{\Sigma}}_{x}{\bm{M}}_{11}+{\bm{M}}_{11}^{\top}{\bm{\Sigma}}_{x}^{\top}\right)\right)}.$ (123)
@@
Building on the linearized-attention experiments, we investigate whether the optimal temperature also benefits more complex Transformer models on linear regression tasks. We evaluate GPT-2 (Radford et al., 2019) under a shift in input covariance (Figure 6). Consistent with prior work (Garg et al., 2022; Zhang et al., 2024), such shifts substantially degrade performance and can even induce nonmonotonic generalization error with respect to context length $l$. Remarkably, applying the optimal temperature mitigates this nonmonotonicity and improves in-context generalization.
@@
Appendix L Rest of the related work
ICL by Transformers —
The ICL capability of Transformers was first brought to prominence by Brown et al. (2020), leading to a surge of empirical and theoretical investigations. Several works have demonstrated that ICL performance improves with model scale Wei et al. (2022); Olsson et al. (2022); Schaeffer et al. (2023), underscoring its importance in modern AI systems. To better understand this phenomenon, synthetic tasks such as linear regression have served as controlled testbeds for analyzing ICL in Transformers (Garg et al., 2022; Zhang et al., 2024; Raventós et al., 2023). A prevailing hypothesis in recent theoretical work is that Transformers implicitly learn algorithms during pretraining, which they subsequently execute during inference (Bai et al., 2023; Li et al., 2023; Akyürek et al., 2023; Ahn et al., 2023; Von Oswald et al., 2023; Mahankali et al., 2024; Fu et al., 2024; Zhang et al., 2024; Li et al., 2024; Park et al., 2024). There remains ongoing debate over the precise nature of these learned procedures. However, our work focuses on a fundamentally different question, which is how attention temperature affects the ICL performance of pretrained Transformers under distribution shifts.

### v2

@@ Appendix F Derivation of the pretraining for ICL by mimicking the Bayes-optimal estimator
Here, we derive the pretraining of the approximate softmax attention model by mimicking the Bayes-optimal ridge estimator (11). The prediction of the approximate softmax attention model can be written as
@@
which makes the prediction of the approximate softmax attention model approximately equal to that of the Bayes-optimal regression. Furthermore, we consider $\tau=1$ for the pretraining. Let’s first focus on $\hat{{\bm{w}}}_{Att}({\bm{C}}_{xx},{\bm{C}}_{xy},C_{yy};{\bm{M}},{\bm{V}})$ as follows
@@
as derived in Appendix B. Looking at equations (60) and (59) together, we can see that setting the parameters as follows would make $\hat{{\bm{w}}}_{Att}=\hat{{\bm{w}}}_{Bayes}$ hold
@@ Appendix G Characterization of generalization error for ICL under distribution shift
by the law of total expectation since ${\bm{w}}_{diff}$ is independent of ${\bm{x}}_{l}$. Note that when writing (72), we safely ignore terms with $(1/l)\bar{{\bm{x}}}_{l}\bar{{\bm{x}}}_{l}^{T}{\bm{v}}_{21}$ in (70) since they vanish as $l\to\infty$ by Assumptions 3.1-3.2 and 4.1. Letting ${\bm{A}}:={\bm{\mu}}_{x}{\bm{\mu}}_{x}^{T}+{\bm{\Sigma}}_{x}$, we write
@@
where we first use the independence of the random variables and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ to simplify the equation. Then, we apply Lemma G.1 and use the fact that $\mathbb{E}[\bar{{\bm{x}}}_{i}\bar{{\bm{x}}}_{i}^{T}]={\bm{\Sigma}}_{x}$ to get the penultimate line. Finally, we drop the vanishing terms and simplify the result using Assumptions 3.1-3.2 and 4.1 in order to reach the last line.
@@
where we again first use the independence of the random variables and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. Then, we apply basic algebraic manipulations. To reach the penultimate line, we utilize Lemma G.1 together with the fact that $\mathbb{E}[\bar{{\bm{x}}}_{i}\bar{{\bm{x}}}_{i}^{T}]={\bm{\Sigma}}_{x}$. Using Assumptions 3.1-3.2 and 4.1, we reach the last line.
@@ Appendix I Analysis of optimal temperature for ICL under distribution shift
$\tau_{\text{opt}}$ $:=\operatorname*{arg\,min}_{\tau}\mathcal{G}(\tau;{\bm{V}},{\bm{M}}),$ (122)
@@
Since $a,b>0$, we reach $\mathcal{G}^{\prime\prime}\left(\frac{2a}{b}\right)=\frac{b^{4}}{8a^{3}}>0$, which means the function has a minimum at $\tau=\frac{2a}{b}$. Therefore, $\tau_{\text{opt}}=\frac{2a}{b}$ is the solution minimizing the generalization error $\mathcal{G}(\tau)$. Writing $a,b$ back into the optimal solution, we get
$\tau_{\text{opt}}=\frac{2\text{Tr}\left({\bm{A}}{\bm{M}}_{11}^{T}{\bm{F}}_{1}{\bm{M}}_{11}\right)}{\text{Tr}\left({\bm{A}}\left({\bm{F}}_{2}{\bm{M}}_{11}+{\bm{M}}_{11}^{T}{\bm{F}}_{2}^{T}\right)\right)},$ (129)
which concludes our derivation of the optimal temperature $\tau_{\text{opt}}$.
@@
In this section, we extract a mathematical heuristic from the optimal temperature in Theorem 4.3 that can be applied to ICL settings beyond our existing setting involving approximate softmax attention and regression tasks. Specifically, we consider Transformers employing standard softmax attention. Recall that the attention temperature scales the pre-softmax scores (i.e., $({\bm{K}}{\bm{Z}})^{\top}({\bm{Q}}{\bm{Z}})$ in (1)), thereby controlling the variance of the final scores. Since the optimal temperature depends on the distribution of these scores, it can be naturally characterized by the moments of that distribution. Our central intuition is that the optimal temperature identified in Theorem 4.3 relates directly to the first two moments of the pre-softmax scores. Although this optimal temperature was derived for approximate softmax attention, the insight remains relevant for softmax attention because the two mechanisms behave similarly in the regime considered (see Appendix D).
@@
$\tau_{\text{opt}}=\frac{v_{22}\operatorname{Tr}\!\left({\bm{\Sigma}}_{x}{\bm{M}}_{11}{\bm{\Sigma}}_{x}{\bm{M}}_{11}^{\top}{\bm{\Sigma}}_{x}\right)}{\frac{1}{2}\operatorname{Tr}\!\left({\bm{\Sigma}}_{x}\left({\bm{\Sigma}}_{x}{\bm{M}}_{11}+{\bm{M}}_{11}^{\top}{\bm{\Sigma}}_{x}^{\top}\right)\right)}.$ (130)
@@
Building on the approximate-attention experiments, we investigate whether the optimal temperature also benefits more complex Transformer models on linear regression tasks. We evaluate GPT-2 (Radford et al., 2019) under a shift in input covariance (Figure 5). Consistent with prior work (Garg et al., 2022; Zhang et al., 2024), such shifts substantially degrade performance and can even induce nonmonotonic generalization error with respect to context length $l$. Remarkably, applying the optimal temperature mitigates this nonmonotonicity and improves in-context generalization.
@@
Appendix L Numerical illustration of the optimal temperature and its generalization behavior in the setting of Section 4.4
To complement the analytical reductions in the setting of Section 4.4, we now present numerical experiments illustrating how the optimal attention temperature varies under different types of distribution shift. These simulations closely follow the structure predicted by the closed-form expression in Theorem 4.3 and its simplified forms in (19) and (22).
Figure 6 shows the optimal temperature as a function of (i) input covariance shift, (ii) task covariance shift, and (iii) noise-variance shift. In each subplot, a single shift parameter ($a$, $b$, or $\sigma$) is varied while the others remain fixed. The closed-form optimal temperatures (16) align closely with the moment-ratio heuristic with correction (22). As anticipated:
  • higher input variance $a$ or noise level $\sigma$ increases the optimal temperature, while
  • task variance $b$ does not significantly change the optimal temperature.
Figure 7 presents the corresponding generalization errors. These results demonstrate that the closed-form characterization accurately captures the key dependencies of the optimal temperature under a range of distribution shifts.
(a) Input Covariance Shifted
(${\bm{\Sigma}}_{x}^{test}=a{\bm{\Sigma}}_{x}^{train})$
(b) Task Covariance Shifted
(${\bm{\Sigma}}_{w}^{test}=b{\bm{\Sigma}}_{w}^{train})$
(c) Noise Shifted
($\sigma^{test}=\sigma)$
Figure 6: Optimal temperature under different types of distribution shift. The moment-ratio heuristic (Appendix J) is derived from the closed-form optimal temperature, with its corrected form given in (22). During training, we use $m=5000$ tasks, noise level $\sigma^{train}=0.1$, means ${\bm{\mu}}_{x}^{train}={\bm{\mu}}_{w}^{train}=\mathbf{0}$, and covariances ${\bm{\Sigma}}_{x}^{train}={\bm{\Sigma}}_{w}^{train}={\bm{I}}$. At test time, we set ${\bm{\mu}}_{x}^{test}={\bm{\mu}}_{w}^{test}=\mathbf{0}$, ${\bm{\Sigma}}_{x}^{test}=aI$, ${\bm{\Sigma}}_{w}^{test}=bI$, and $\sigma^{test}=\sigma$. In each subplot, exactly one of $a$, $b$, or $\sigma$ is varied, while the other two remain fixed at their training-distribution values to isolate the effect of a single shift dimension. The dimension and context length are set to $d=50$ and $l=2d$.
(a) Input Covariance Shifted
(${\bm{\Sigma}}_{x}^{test}=a{\bm{\Sigma}}_{x}^{train})$
(b) Task Covariance Shifted
(${\bm{\Sigma}}_{w}^{test}=b{\bm{\Sigma}}_{w}^{train})$
(c) Noise Shifted
($\sigma^{test}=\sigma)$
Figure 7: Generalization errors corresponding to Figure 6.